# MAFA: A multi-agent framework for annotation

### v1

@@ Abstract
Modern applications require accurate and efficient retrieval of information in response to user queries. Mapping user utterances to the most relevant Frequently Asked Questions (FAQs) is a crucial component of these systems. Traditional approaches often rely on a single model or technique, which may not capture the nuances of diverse user inquiries. In this paper, we introduce a multi-agent framework for FAQ annotation that combines multiple specialized agents with different approaches and a judge agent that reranks candidates to produce optimal results. Our agents utilize a structured reasoning approach inspired by Attentive Reasoning Queries (ARQs), which guides them through systematic reasoning steps using targeted, task-specific JSON queries. Our framework features a specialized few-shot example strategy, where each agent receives different few-shots, enhancing ensemble diversity and coverage of the query space. We evaluate our framework on a real-world banking dataset as well as public benchmark datasets (LCQMC and FiQA), demonstrating significant improvements over single-agent approaches across multiple metrics, including a 14% increase in Top-1 accuracy, an 18% increase in Top-5 accuracy, and a 12% improvement in Mean Reciprocal Rank on our dataset, and similar gains on public benchmarks when compared with traditional single agent annotation techniques. Our framework is particularly effective at handling ambiguous queries, making it well-suited for deployment in production applications while showing strong generalization capabilities across different domains and languages.
 [^1] JP Morgan Chase CCB ML Team
@@ 1 Introduction
In this paper, we propose a multi-agent framework for FAQ annotation (MAFA). Our framework employs specialized agents to generate candidate FAQs, followed by a judge agent that reranks these candidates to produce a final, optimized ranking. This approach is inspired by recent advances in ensemble learning and LLM-based multi-agent systems introduced by Du et al. (2023).
A key innovation in our framework is the use of specialized few-shot examples for each agent. Rather than providing identical examples to all agents, we carefully select different sets that align with each agent’s strengths and capabilities, maximizing effectiveness on particular query types while increasing the overall diversity of the system’s coverage. This novel multi-agent framework for FAQ annotation combines diverse approaches with a dedicated judge agent that effectively reranks candidate FAQs, resulting in significant performance improvements. Our comprehensive evaluation on both a real-world banking dataset and public benchmarks demonstrates the framework’s effectiveness, with particular analysis of performance across different query types. The implementation reveals practical insights for deploying such systems in production banking environments, addressing challenges like ambiguous queries and technical terminology. Together, these contributions establish a robust methodology for improving FAQ annotation accuracy while maintaining efficiency in customer-facing applications.
@@ 2 Related work
Figure 1: Architecture of the MAFA framework. Each agent receives diversified few-shot examples, generating candidate FAQs that are then reranked by the judge agent.
@@ 2.1 FAQ retrieval systems
Retrieval has been a focus of research for decades, evolving from simple keyword matching to sophisticated neural approaches. Early systems relied on information retrieval techniques like TF-IDF and BM25 (Robertson and Zaragoza (2009)). With the advent of neural networks, embedding-based approaches became prevalent, including sentence-BERT (Reimers and Gurevych (2019)) and other transformer-based models that encode queries and FAQs into a shared semantic space.
@@ 2.2 Multi-agent systems
Multi-agent systems have gained popularity with the rise of powerful LLMs. These systems typically involve multiple agents with specialized roles collaborating to solve complex reasoning tasks as showed by Du et al. (2023); Hegazy (2024). Park et al. (2023) demonstrated the effectiveness of generative agents for collaborative problem-solving, while similarly, Wu and others (2023); Li et al. (2023) presented approaches for creating autonomous agents that can collaborate through natural language communication.
These works lay the groundwork for our multi-agent FAQ annotation system, though they did not specifically address the information retrieval domain. Our work draws inspiration from these approaches but focuses specifically on the FAQ annotation task in the banking domain, where precision and domain knowledge are crucial.
@@ 2.3 Retrieval-augmented generation (RAG)
Retrieval-Augmented Generation (RAG) has emerged as a powerful paradigm for enhancing large language models with external knowledge. Since its introduction by Lewis and others (2020), numerous advances have been made in improving retrieval quality, reranking strategies, and generation capabilities. Recent work has focused on enhancing RAG with more sophisticated retrieval mechanisms Lewis and others (2020), hybrid search strategies Zhao et al. (2024); Kim et al. (2025), and context-aware generation Su et al. (2023).
The original RAG framework presented by Lewis and others (2020) introduced the key concept of combining neural retrievers with sequence-to-sequence models for knowledge-intensive NLP tasks. This approach allowed models to access non-parametric knowledge without expanding model size. Extensions by Izacard et al. (2023) demonstrated that RAG systems can be adapted to few-shot learning scenarios, while Borgeaud et al. (2022) introduced more efficient retrieval mechanisms for large-scale knowledge bases.
@@ 2.4 Judging and reranking in information retrieval
Reranking has been widely used in information retrieval to improve the relevance of search results as shown by Liu (2009). Recent approaches leverage LLMs for reranking, as demonstrated by Nogueira et al. (2020), who used BERT for document reranking. Our approach extends this concept to FAQ annotation, with a specialized judge agent that considers multiple sources of evidence.
Recent work by Niu et al. (2025) introduced JudgeRank, a framework that leverages large language models for reasoning-intensive reranking tasks. This approach demonstrated the potential of using LLMs as judges to evaluate the quality of retrieved content, a concept we extend in our work. Similarly, Wang et al. (2023) showed how self-consistency can improve the reliability of LLM judgments, a principle we incorporate into our judge agent design.
@@ 2.5 Structured reasoning for large language models
Recent research has shown that structured reasoning approaches can significantly improve LLM performance in complex tasks. Karov et al. (2025) introduced Attentive Reasoning Queries (ARQs), which guide LLMs through systematic reasoning steps using targeted, task-specific queries in a structured JSON format. This approach has been shown to outperform free-form Chain-of-Thought reasoning by directing the model’s attention to critical instructions and facilitating intermediate reasoning steps. ARQs counteract the ’lost in the middle’ phenomenon in autoregressive models, where information in the middle of long prompts tends to receive less attention than content at the beginning or end as shown by Liu et al. (2024). By structuring the reasoning process with explicit queries in a JSON schema, ARQs ensure that important information remains in the model’s focus when it’s most needed for reasoning and decision-making. This approach has proven particularly effective for tasks requiring strict adherence to guidelines and prevention of hallucinations, which aligns well with our requirements for FAQ annotation.
@@ 4.1 Query planning agent
Powered by OpenAI et al. (2024)’s GPT-4o, the agent follows Kim et al. (2025)’s approach that combines lexical expansion with semantic understanding. Our implementation enhances this methodology with intent-based expansion, focusing specifically on terms related to the inferred user intent, thereby improving relevance and retrieval performance.
@@ 4.2 Ranker agent network
We deploy a complementary ensemble of four specialized agents to generate candidate FAQs, each leveraging distinct information retrieval paradigms. The FAQ Mapper without embeddings employs sophisticated prompt engineering with detailed instructions for deep intent analysis and banking-specific considerations, generating structured outputs with precise intent categorization and explicit reasoning chains. Building on this foundation, the FAQ Mapper with Embeddings combines dense vector representation techniques from Karpukhin et al. (2020) with our enhanced prompting strategy, enabling focused semantic retrieval and ranking of a curated candidate set. Two additional agents incorporate FAQ answer content: one without embeddings for contextual reasoning, and another implementing a hybrid approach that fuses embedding-based retrieval with comprehensive answer-aware evaluation, drawing on innovations in retrieval-augmented generation described by Lewis and others (2020). This multi-faceted approach ensures robust performance across diverse query types and varying levels of query ambiguity.
@@ 4.2.1 Structured agent prompting
The design of our agent prompts is inspired by the ARQs approach introduced by Karov et al. (2025). Rather than using free-form Chain-of-Thought (CoT) reasoning, we employ a structured JSON format with targeted queries that guide each agent through systematic reasoning steps. This approach offers several advantages over traditional reasoning methods.
First, the JSON structure allows us to explicitly remind the LLM of key instructions at critical decision points, mitigating the "lost in the middle" phenomenon, explored by Liu et al. (2024), where important information receives less attention. Second, by breaking down the reasoning process into specific queries, we can ensure that each step in the agent’s decision-making is explicit and traceable, facilitating intermediate reasoning. Third, the structured format makes it easier to inspect and debug the reasoning process, as each component of the decision is clearly demarcated. Finally, unlike general-purpose CoT, our structured queries incorporate domain knowledge about banking and FAQ annotation to address task-specific challenges.
@@
Each ranker agent in our system produces a reordered list of candidates with carefully calibrated scores that reflect their specific evaluation criteria. These independently ranked lists are then passed to the judge agent for final evaluation and synthesis. Our implementation draws on established research in information retrieval and ranking techniques. The semantic ranker builds on the cross-encoder approach proposed by Nogueira et al. (2020), which has demonstrated strong performance for ranking in information retrieval tasks. For the intent ranker, we took inspiration from the intent-aware ranking model developed by Su et al. (2023), while our relevance ranker incorporates multiple assessment factors following the comprehensive approach outlined by Niu et al. (2025).
@@ 4.3 Judge agent
Our judge agent implements the "judge as a judge" mechanism introduced by Liu et al. (2025), where multiple judgments are generated and then meta-evaluated to select the most reliable one. This approach significantly improves the stability and accuracy of the final rankings.
The consistency verification mechanism is also inspired by the work of Liu et al. (2025), which demonstrated the benefits of consistency-based evaluation for language model judgments. Our multi-dimensional evaluation extends this approach by considering multiple aspects of relevance, similar to the approach of Wang et al. (2023).
@@ 4.4 Implementation
Our implementation leverages a strategic combination of models: OpenAI’s text-embedding-ada-002 for semantic representation and GPT-4o as the base LLM for our agents.
@@
For each agent we randomly sample without replacement unique examples and ensure each agent receives exactly five. The use of different few-shot examples is theoretically grounded in the principle of ensemble diversity. By providing each agent with different examples, we increase the diversity of their outputs, which is known to improve ensemble performance as studied by Dietterich (2000); Kuncheva and Whitaker (2003). This approach also aligns with the concept of mixture of experts as introduced by Jacobs et al. (1991), where different models specialize in different parts of the input space.
In the context of LLM few-shot learning, specialized examples can be seen as a form of prompt engineering that tailors each agent’s behavior to its strengths. Research has shown that carefully selected few-shot examples can significantly impact LLM performance on specific tasks as shown by Brown and others (2020); Liu et al. (2023).
@@
As described in Section 4.2.1, all our agents use structured JSON-based prompting inspired by the ARQ framework Karov et al. (2025). This approach guides the LLM through a systematic reasoning process with targeted queries that focus attention on critical aspects of the task.
@@
The judge agent serves as the final arbiter in our framework, reranking all candidates to produce an optimized ranking. It processes the original query, candidate FAQs with their scores and reasoning, agent recommendation details, training examples, and complete FAQ content. This comprehensive input allows the judge to reevaluate each candidate, assign calibrated relevance scores, and provide detailed reasoning that enhances system interpretability and auditability.
@@
4:     for each $agent$ in $specialized\_agents$ do
@@
7:         $all\_candidates.extend(candidates)$
@@
11:     return $reranked\_candidates[:5]$triangleright Return top 5
@@
Embedding generation
We use the OpenAI text-embedding-ada-002 model to generate embeddings for FAQs. For agents that consider answer context, we concatenate the question and answer text before generating embeddings, using the format: "Question: {question} Answer: {answer}".
4.4.1 LLM inference
We use the OpenAI GPT-4o model for all LLM inference tasks. For specialized agents, we use a temperature of 0.1 to ensure consistent outputs, while for the judge agent, we use a slightly higher temperature of 0.3 to allow for more nuanced reasoning.
@@ 4.4.1 LLM inference
Caching and fallback mechanisms
In case of API failures or other issues, we implement fallback mechanisms that ensure the system can still generate predictions. If the judge agent fails, we fall back to a simple aggregation approach that ranks FAQs based on their average scores across all agents. We also implement caching to avoid redundant API calls for repeated utterances.
@@
The judge agent outputs a structured JSON response containing reranked FAQs with titles, relevance scores (0-100), and detailed reasoning explaining each ranking decision. This format ensures consistency and facilitates both human review and automated processing of results.
@@ 5 Experimental setup
  • Bank dataset: A real-world major bank application user utterances dataset, collected over 1 year, with human expert-annotated ranked FAQs serving as our ground-truth labels.
  • LCQMC dataset: The Large-scale Chinese Question Matching Corpus Liu et al. (2018), which contains pairs of questions with labels indicating whether they have the same intent.
  • FiQA dataset: A financial domain question answering dataset from the Financial Question Answering Task of the WWW’18 conference Maia et al. (2018), consisting of user questions about financial topics with expert-provided answers.
@@
For the public benchmarks:
@@
The LCQMC dataset by Liu et al. (2018) is designed for evaluating question matching capabilities, which aligns with our goal of mapping user utterances to FAQs. It contains pairs of questions with binary labels indicating whether they have the same intent. We converted this into an FAQ annotation task by treating the second question in each pair as an FAQ question.
The FiQA dataset by Maia et al. (2018) contains financial domain questions paired with expert answers, making it particularly relevant for our banking application scenario. We processed it to create an FAQ annotation task by treating the original questions as FAQ questions and generating variations of these questions as user utterances.

### v3

@@ Abstract
Modern consumer banking applications require accurate and efficient retrieval of information in response to user queries. Mapping user utterances to the most relevant Frequently Asked Questions (FAQs) is a crucial component of these systems. Traditional approaches often rely on a single model or technique, which may not capture the nuances of diverse user inquiries. In this paper, we introduce a multi-agent framework for FAQ annotation that combines multiple specialized agents with different approaches and a judge agent that reranks candidates to produce optimal results. Our agents utilize a structured reasoning approach inspired by Attentive Reasoning Queries (ARQs), which guides them through systematic reasoning steps using targeted, task-specific JSON queries. Our framework features a few-shot example strategy, where each agent receives different few-shots, enhancing ensemble diversity and coverage of the query space. We evaluate our framework on a real-world major bank dataset as well as public benchmark datasets (LCQMC and FiQA), demonstrating significant improvements over single-agent approaches across multiple metrics, including a 14% increase in Top-1 accuracy, an 18% increase in Top-5 accuracy, and a 12% improvement in Mean Reciprocal Rank on our dataset, and similar gains on public benchmarks when compared with traditional and single-agent annotation techniques. Our framework is particularly effective at handling ambiguous queries, making it well-suited for deployment in production banking applications while showing strong generalization capabilities across different domains and languages.
[…]
123
@@ 1 Introduction
In this paper, we propose a multi-agent framework for FAQ annotation (MAFA). Our framework employs specialized agents to generate candidate FAQs, followed by a judge agent that reranks these candidates to produce a final, optimized ranking. This approach is inspired by recent advances in ensemble learning and LLM-based multi-agent systems introduced by Du et al. [3].
Our framework achieves significant performance gains through strategic few-shot example selection tailored to individual agents. Rather than applying uniform examples across all agents, we curate distinct example sets that align with each agent’s unique capabilities and strengths. This targeted approach enhances performance across diverse query types while expanding the system’s coverage and adaptability. The multi-agent architecture, combined with a dedicated judge agent for intelligent reranking, delivers substantial improvements in FAQ mapping accuracy. Through comprehensive evaluation on both proprietary banking data and public benchmarks (LCQMC and FiQA), we demonstrate the framework’s effectiveness across varied domains and query patterns. Our implementation provides valuable insights for production deployment in banking environments, effectively handling challenges such as ambiguous queries and domain-specific terminology. These contributions establish a robust and practical methodology that advances FAQ annotation accuracy while maintaining the efficiency required for customer-facing applications.
@@ 2.1 FAQ retrieval systems
Retrieval has been a focus of research for decades, evolving from simple keyword matching to sophisticated neural approaches. Early systems relied on information retrieval techniques like TF-IDF and BM25 [24]. With the advent of neural networks, embedding-based approaches became prevalent, including sentence-BERT [23] and other transformer models that encode queries and FAQs into a shared semantic space.
@@ 2.2 Multi-agent systems
Multi-agent systems have gained popularity with the rise of powerful LLMs. These systems typically involve multiple agents with specialized roles collaborating to solve complex reasoning tasks as showed by [3, 4]. Park et al. [22] demonstrated the effectiveness of generative agents for collaborative problem-solving, while similarly, Wu et al. [27], Li et al. [12] presented approaches for creating autonomous agents that can collaborate through natural language communication. These works lay the groundwork for our multi-agent FAQ annotation system, though they did not specifically address the information retrieval domain. Our work draws inspiration from these approaches but focuses specifically on the FAQ annotation task in the banking domain, where precision and domain knowledge are crucial.
@@ 2.3 Retrieval-augmented generation (RAG)
Retrieval-Augmented Generation (RAG) has emerged as a powerful paradigm for enhancing large language models with external knowledge. Since its introduction by Lewis et al. [11], numerous advances have been made in improving retrieval quality, reranking strategies, and generation capabilities. Recent work by Lewis et al. [11] has focused on enhancing RAG with more sophisticated retrieval mechanisms, hybrid search strategies [28, 9], and context-aware generation [25].
@@ 2.4 Judging and reranking in information retrieval
Reranking has been widely used in information retrieval to improve the relevance of search results as shown by Liu [16]. Recent approaches leverage LLMs for reranking, as demonstrated by Nogueira et al. [20], who used BERT for document reranking. Our approach extends this concept to FAQ annotation, with a specialized judge agent that considers multiple sources of evidence.
Recent work by Niu et al. [19] introduced JudgeRank, a framework that leverages large language models for reasoning-intensive reranking tasks. This approach demonstrated the potential of using LLMs as judges to evaluate the quality of retrieved content, a concept we extend in our work. Similarly, Wang et al. [26] showed how self-consistency can improve the reliability of LLM judgments, a principle we incorporate into our judge agent design.
@@ 2.5 Structured reasoning for large language models
Recent research has shown that structured reasoning approaches can significantly improve LLM performance in complex tasks. Karov et al. [7] introduced Attentive Reasoning Queries (ARQs), which guide LLMs through systematic reasoning steps using targeted, task-specific queries in a structured JSON format. This approach has been shown to outperform free-form Chain-of-Thought reasoning by directing the model’s attention to critical instructions and facilitating intermediate reasoning steps. ARQs counteract the ’lost in the middle’ phenomenon in autoregressive models, where information in the middle of long prompts tends to receive less attention than content at the beginning or end as shown by Liu et al. [13]. By structuring the reasoning process with explicit queries in a JSON schema, ARQs ensure that important information remains in the model’s focus when it’s most needed for reasoning and decision-making. This approach has proven particularly effective for tasks requiring strict adherence to guidelines and prevention of hallucinations, which aligns well with our requirements for FAQ annotation.
@@ 3 Problem formulation
Figure 1: Architecture of MAFA. Each agent generates candidate FAQ annotations that are re-ranked by the judge agent. All agents have access to the FAQ database.
@@ 4.1 Query planning agent
Powered by OpenAI et al. [21]’s GPT-4o, the agent follows Kim et al. [9]’s approach that combines lexical expansion with semantic understanding. Our implementation enhances this methodology with intent-based expansion, focusing specifically on terms related to the inferred user intent, thereby improving relevance and retrieval performance.
@@ 4.2 Ranker agent network
We deploy a complementary ensemble of four specialized agents to generate candidate FAQs, each leveraging distinct information retrieval paradigms. The FAQ Mapper without embeddings employs sophisticated prompt engineering with detailed instructions for deep intent analysis and banking-specific considerations, generating structured outputs with precise intent categorization and explicit reasoning chains. Building on this foundation, the FAQ Mapper with Embeddings combines dense vector representation techniques from Karpukhin et al. [8] with our enhanced prompting strategy, enabling focused semantic retrieval and ranking of a curated candidate set. Two additional agents incorporate FAQ answer content: one without embeddings for contextual reasoning, and another implementing a hybrid approach that fuses embedding-based retrieval with comprehensive answer-aware evaluation, drawing on innovations in retrieval-augmented generation described by Lewis et al. [11]. This multi-faceted approach ensures robust performance across diverse query types and varying levels of query ambiguity.
@@ 4.2.1 Structured agent prompting
The design of our agent prompts is inspired by the ARQs approach introduced by Karov et al. [7]. Rather than using free-form Chain-of-Thought (CoT) reasoning, we employ a structured JSON format with targeted queries that guide each agent through systematic reasoning steps. This approach offers several advantages over traditional reasoning methods.
First, the JSON structure allows us to explicitly remind the LLM of key instructions at critical decision points, mitigating the "lost in the middle" phenomenon, explored by Liu et al. [13], where important information receives less attention. Second, by breaking down the reasoning process into specific queries, we can ensure that each step in the agent’s decision-making is explicit and traceable, facilitating intermediate reasoning. Third, the structured format makes it easier to inspect and debug the reasoning process, as each component of the decision is clearly demarcated. Finally, unlike general-purpose CoT, our structured queries incorporate domain knowledge about banking and FAQ annotation to address task-specific challenges.
@@
Each ranker agent in our system produces a reordered list of candidates with carefully calibrated scores that reflect their specific evaluation criteria. These independently ranked lists are then passed to the judge agent for final evaluation and synthesis. Our implementation draws on established research in information retrieval and ranking techniques. The semantic ranker builds on the cross-encoder approach proposed by Nogueira et al. [20], which has demonstrated strong performance for ranking in information retrieval tasks. For the intent ranker, we took inspiration from the intent-aware ranking model developed by Su et al. [25], while our relevance ranker incorporates multiple assessment factors following the comprehensive approach outlined by Niu et al. [19].
@@ 4.3 Judge agent
Our judge agent implements the "judge as a judge" mechanism introduced by Liu et al. [15], where multiple judgments are generated and then meta-evaluated to select the most reliable one. This approach significantly improves the stability and accuracy of the final rankings.
The consistency verification mechanism is also inspired by the work of [15], which demonstrated the benefits of consistency-based evaluation for language model judgments. Our multi-dimensional evaluation extends this approach by considering multiple aspects of relevance, similar to the approach of Wang et al. [26].
@@ 4.4 Implementation
Although we use the OpenAI GPT-4o model for all LLM inference tasks, our framework allows for easy adaptation to other LLM families such as Claude, Gemini or Llama. For ranker agents, we use a temperature of 0.1 to ensure consistent outputs, while for the judge agent, we use a slightly higher temperature of 0.3 to allow for more nuanced reasoning.
Framework implementation
While we developed a custom orchestration system for research flexibility, our architecture is compatible with established multi-agent frameworks. The system can be readily implemented using LangChain’s agent executors or OpenAI’s Assistants API. For production deployment, we recommend using these established frameworks to benefit from their robust error handling, monitoring, and scaling capabilities. Our implementation abstracts the agent coordination logic, making it framework-agnostic and portable.
@@
For each agent we randomly sample without replacement unique examples and ensure each agent receives exactly five. The use of different few-shot examples is theoretically grounded in the principle of ensemble diversity. By providing each agent with different examples, we increase the diversity of their outputs, which is known to improve ensemble performance as studied by Dietterich [2], Kuncheva and Whitaker [10]. This approach also aligns with the concept of mixture of experts [6], where different models specialize in different parts of the input space.
In the context of LLM few-shot learning, specialized examples can be seen as a form of prompt engineering that tailors each agent’s behavior to its strengths. Research has shown that carefully selected few-shot examples can significantly impact LLM performance on specific tasks [1, 14].
@@
As described in Section 4.2.1, all our agents use structured JSON-based prompting inspired by the ARQ framework showcased by Karov et al. [7]. This approach guides the LLM through a systematic reasoning process with targeted queries that focus attention on critical aspects of the task.
Embedding retrieval in ranker agents
Two of the four ranker agents utilize embedding retrieval and Approximate Nearest Neighbors (ANN) [5] to identify the 50 most relevant FAQ neighbors to the user’s query, effectively narrowing the search space for annotation. We employ the OpenAI text-embedding-large model to generate the embeddings. For agents that take answer context into account, we concatenate the question and answer text before generating embeddings, using the format: "Question: question Answer: answer".
@@
The judge agent serves as the final arbiter in our framework, reranking all single-agent candidates to produce an optimized ranking. It processes the original query, candidate FAQs with their scores and reasoning, agent recommendation details, training examples, and complete FAQ content. This comprehensive input allows the judge to reevaluate each candidate, assign calibrated relevance scores, and provide detailed reasoning that enhances system interpretability and auditability.
@@
4:  for each agentagent in specialized_agentsspecialized\_agents do
@@
7:   all_candidates.extend(candidates)all\_candidates.extend(candidates)
@@
11:  return reranked_candidates[:5]reranked\_candidates[:5] $\triangleright$ Return top 5
@@
The judge agent outputs a structured JSON response containing reranked FAQs with titles, relevance scores (0-100), detailed reasoning explaining each ranking decision, and confidence in mapping (HIGH/MED/LOW). This format ensures consistency and facilitates both human review and automated processing of results.
@@ 5.1 Datasets
  • Bank dataset: A real-world major bank application user utterances dataset, collected over 1 year in production, with human expert-annotated ranked FAQs serving as our ground-truth labels.
  • LCQMC dataset: The Large-scale Chinese Question Matching Corpus [17] comprises pairs of questions labeled to indicate whether they share the same intent. This corpus was chosen to showcase the effectiveness of our framework with multilingual FAQs that are not related to finance.
  • FiQA dataset: A financial domain question answering dataset from the Financial Question Answering Task of the WWW’18 conference [18], consisting of user questions about financial topics with expert-provided answers.
@@
Likewise for the public benchmarks table 2:
@@
The LCQMC dataset by Liu et al. [17] is designed for evaluating question matching capabilities, which aligns with our goal of mapping user utterances to FAQs. It contains pairs of questions with binary labels indicating whether they have the same intent. We converted this into an FAQ annotation task by treating the second question in each pair as an FAQ question.
The FiQA dataset by Maia et al. [18] contains financial domain questions paired with expert answers, making it particularly relevant for our banking application scenario. We processed it to create an FAQ annotation task by treating the original questions as FAQ questions and generating variations of these questions as user utterances.